OpenStudy (anonymous):

I need some clarification with derivatives please. What would the derivative of 4Sec(2x)^3 be???

OpenStudy (anonymous):

\[4Sec ^{3}(2x)\]

hartnn (hartnn):

u need clarification mean you have tried, what u got ?

OpenStudy (anonymous):

OK, I know im suppoased to bering down the three. so the first part is 12. and the derivative of sec(x) is Sec(x)Tan(x). So the second part is Sec(2x)Tan(2x). But where the the power of 2 go? So far I have 12Sec(2x)Tan(2x) I just dont' know where the "squared" goes....

OpenStudy (anonymous):

use the chain rule, it will be the derivative of the outside times the derivative of the inside

OpenStudy (anonymous):

12Sec(2x)Tan(2x)*2

OpenStudy (anonymous):

what is the derivative of the outside?

OpenStudy (anonymous):

it will be just, 12(sec(2x))^2 right?

hartnn (hartnn):

the chain rule says f(g(x))' = f'(g(x))g'(x) so here first derivate this according to rule x^3 ,(3x^2) so, what u get here ?

OpenStudy (anonymous):

It's actualy (sec(2x))^3 I wrote it wrong the first time. does that change anything?

hartnn (hartnn):

no.

OpenStudy (anonymous):

um... I'm not undertanding what you mean ._.

hartnn (hartnn):

applied chain rule before ?

OpenStudy (anonymous):

Yeah, the only thing that is throwing me off is the "sec"

hartnn (hartnn):

ok, i will write 1st step , see whether it make sense. d/dx [4sec^3 (2x)] = 4*3 sec^2 (2x) d/dx[sec(2x)]

hartnn (hartnn):

i considered sec^3 2x = u^3 then its derivative will be 3u^2 du/dx

OpenStudy (anonymous):

wait, so sec does't turn into sec(x)tan(x)??

hartnn (hartnn):

it will... what is d/dx [sec (2x)] ??

OpenStudy (anonymous):

Sec(2x)Tan(2x)2?

hartnn (hartnn):

yes, correct ! so just put that in 4*3 sec^2 (2x) d/dx[sec(2x)]

hartnn (hartnn):

did u get this step ? 4*3 sec^2 (2x) d/dx[sec(2x)]

OpenStudy (anonymous):

yes, 12Sec^2(2x)Tan^2(2x)2?

hartnn (hartnn):

won't that be sec^3 (2x) ?

hartnn (hartnn):

4*3 sec^2 (2x) . Sec(2x)Tan(2x)*2 = 24 sec^3 (2x) tan (2x) got this ?

OpenStudy (anonymous):

Ok, I think I get it. But why does it remain as 3rd power? I thought it was suppoased to be one less?

hartnn (hartnn):

because one extra power got added when u diff. sec 2x

OpenStudy (anonymous):

I don't see it 0_0

hartnn (hartnn):

2 powers form sec^3 2x---->3sec^2 (2x) and one power from sec(2x) ----> sec2x tan 2x so what is 3sec^2 (2x) * (sec2x tan 2x) = ?

OpenStudy (anonymous):

3Sec^3(2x)Tan(2x)?

hartnn (hartnn):

right, now u see how sec^3 2x comes ? and for linear function, derivative is always one power less d/du (u^n) = n u^(n-1) sec x is not a linear funtion

hartnn (hartnn):

any confusion still ?

OpenStudy (anonymous):

Sorry, my computer died. But yes! I get it! Thank you.